Consider the decomposition of a metal oxide to its elements, where M represents a generic metal.

M2O3(s)−⇀↽− 2M(s)+32O2(g)


Substance
ΔG°f(kJ/mol)
M2O3(s) −6.90
M(s) 0
O2(g) 0

What is the standard change in Gibbs energy for the reaction, as written, in the forward direction?

ΔG∘rxn=

kJ/mol

What is the equilibrium constant of this reaction, as written, in the forward direction at 298 K?

K=

What is the equilibrium pressure of O2(g) over M(s) at 298 K?

PO2=
Answer #1

Mooiz ere +39(3) D Goxh = ({ shf) - (ESGP ) Product > Reactant = (2 * 0 -3xo)-( - 6 -3) = 0.9 kJ/mol Ang Go = -RT luke 6900 =